### v1

@@ Appendix A Poloidal-toroidal decomposition of the magnetic field
$\vec{B}_{\rm pol}=\vec{\nabla}\times(\vec{\nabla}\times\Phi\vec{k})~,$ (78)
$\vec{B}_{\rm tor}=\vec{\nabla}\times\Psi\vec{k}~,$ (79)
@@
$\vec{B}_{\rm pol}=\vec{\nabla}\times(\vec{\nabla}\Phi\times\vec{r})=-\vec{r}~\nabla^{2}\Phi+\vec{\nabla}\left(\frac{\partial(r\Phi)}{\partial r}\right)~,$ (80)
@@
Generally speaking, the radial component of the magnetic field is included in the poloidal part, while the $\theta$ and $\varphi$ components are shared between poloidal and toroidal components. In axial symmetry, $\Phi=\Phi(r,\theta)$ and $\Psi=\Psi(r,\theta)$, the expressions are further simplified: the toroidal magnetic field is directed along the azimuthal direction $\hat{\varphi}$. In this case the potential vector is purely azimuthal and given by $\vec{A_{\varphi}}=-\vec{r}\times\vec{\nabla}\Phi~,$ and the poloidal field can be directly derived from $\vec{B}_{\rm pol}=\vec{\nabla}\times\vec{A_{\varphi}}~.$
@@
$P(r,\theta)=r\sin\theta~A_{\varphi}(r,\theta)~,$ (83)
@@
$\vec{B}_{\rm pol}=\frac{\vec{\nabla}P(r,\theta)\times\hat{\varphi}}{r\sin\theta}~,$ (84)
$\vec{B}_{\rm tor}=(\vec{\nabla}\xi)_{\rm pol}\times(\vec{\nabla}P)_{\rm pol}\equiv\frac{T}{r\sin\theta}\hat{\varphi}~$, (85)
where we have introduced the scalar stream function $T$ used, for instance, in Akgün et al (2017) and following works (in the force-free case, $T$ is a function of $P$, see Sect. 5.2). The conversion between the two formalisms in axial symmetry is shown in Table LABEL:tab:formalism.
Formalisms Akgün et al (2017) Kojima (2017) Geppert and Wiebicke (1991)
@@
Table 1: Comparison between different notations in axial symmetry. Pons et al (2009b) used the same notation as Geppert and Wiebicke (1991), and in Gourgouliatos et al (2016) $\Phi$ and $\Psi$ are denominated by $V_{p}$ and $V_{t}$, respectively.
@@ Appendix B Potential solutions with Green’s method
$2\pi\chi_{m}(\vec{r})=-\int_{S}\frac{\partial G}{\partial n^{\prime}}(\vec{r},\vec{r}^{\prime})\chi_{m}(\vec{r}^{\prime}){\rm d}S^{\prime}+\int_{S}G(\vec{r},\vec{r}^{\prime})\frac{\partial\chi_{m}}{\partial n^{\prime}}(\vec{r}^{\prime}){\rm d}S^{\prime}~,$ (88)
@@
$B_{\theta}(R,\theta)\equiv\frac{1}{R}\frac{\partial\chi_{m}}{\partial\theta}(R,\theta)~,$ (90)
@@
$G(R,\theta,\theta^{\prime},\varphi^{\prime})=\frac{1}{\sqrt{2}R}\left[1-\cos(\theta-\theta^{\prime})+2\sin\theta\sin\theta^{\prime}\sin^{2}\left(\frac{\varphi^{\prime}}{2}\right)\right]^{-1/2}~,$ (94)

### v2

@@ Acknowledgements.
Declarations
Conflict of interest The authors declare no conflict of interest.
@@ Appendix A Poloidal-toroidal decomposition of the magnetic field
$\vec{B}_{\rm pol}=\vec{\nabla}\times(\vec{\nabla}\times\Phi\vec{k})\,$, (78)
$\vec{B}_{\rm tor}=\vec{\nabla}\times\Psi\vec{k}\,$, (79)
@@
$\vec{B}_{\rm pol}=\vec{\nabla}\times(\vec{\nabla}\Phi\times\vec{r})=-\vec{r}~\nabla^{2}\Phi+\vec{\nabla}\left(\frac{\partial(r\Phi)}{\partial r}\right)\,$, (80)
@@
Generally speaking, the radial component of the magnetic field is included in the poloidal part, while the $\theta$ and $\varphi$ components are shared between poloidal and toroidal components. In axial symmetry, $\Phi=\Phi(r,\theta)$ and $\Psi=\Psi(r,\theta)$, the expressions are further simplified: the toroidal magnetic field is directed along the azimuthal direction $\hat{\varphi}$. In this case the potential vector is purely azimuthal and given by $\vec{A_{\varphi}}=-\vec{r}\times\vec{\nabla}\Phi\,$, and the poloidal field can be directly derived from $\vec{B}_{\rm pol}=\vec{\nabla}\times\vec{A_{\varphi}}~.$
@@
$P(r,\theta)=r\sin\theta~A_{\varphi}(r,\theta)\,$, (83)
@@
$\vec{B}_{\rm pol}=\frac{\vec{\nabla}P(r,\theta)\times\hat{\varphi}}{r\sin\theta}\,$, (84)
$\vec{B}_{\rm tor}=(\vec{\nabla}\xi)_{\rm pol}\times(\vec{\nabla}P)_{\rm pol}\equiv\frac{T}{r\sin\theta}\hat{\varphi}\,$, (85)
where we have introduced the scalar stream function $T$ used, for instance, in Akgün et al. (2017) and following works (in the force-free case, $T$ is a function of $P$, see Sect. 5.2). The conversion between the two formalisms in axial symmetry is shown in Table LABEL:tab:formalism.
Table 1: Comparison between different notations in axial symmetry. Pons et al. (2009) used the same notation as Geppert and Wiebicke (1991), and in Gourgouliatos et al. (2016) $\Phi$ and $\Psi$ are denominated by $V_{p}$ and $V_{t}$, respectively.
Formalisms Akgün et al. (2017) Kojima (2017) Geppert and Wiebicke (1991)
@@ Appendix B Potential solutions with Green’s method
$2\pi\chi_{m}(\vec{r})=-\int_{S}\frac{\partial G}{\partial n^{\prime}}(\vec{r},\vec{r}^{\prime})\chi_{m}(\vec{r}^{\prime}){\rm d}S^{\prime}+\int_{S}G(\vec{r},\vec{r}^{\prime})\frac{\partial\chi_{m}}{\partial n^{\prime}}(\vec{r}^{\prime}){\rm d}S^{\prime}\,$, (88)
@@
$B_{\theta}(R,\theta)\equiv\frac{1}{R}\frac{\partial\chi_{m}}{\partial\theta}(R,\theta)\,$, (90)
@@
$G(R,\theta,\theta^{\prime},\varphi^{\prime})=\frac{1}{\sqrt{2}R}\left[1-\cos(\theta-\theta^{\prime})+2\sin\theta\sin\theta^{\prime}\sin^{2}\left(\frac{\varphi^{\prime}}{2}\right)\right]^{-1/2}\,$, (94)